Could you please,help me solve
n/2-1+n/5=-1/2

Answer by tutorcecilia(2144)

10(n/2-1+n/5=-1/2) [Multiply each term by a 10]
5n-10+2n=-5 [Simplify]
7n-10=-5 [Solve for n]
7x=-5+10
7x=5
x=5/7
.
Check by plugging the value of x back into the original equation]